# Blind Source Separation of Single-Channel Mixtures via Multi-Encoder Autoencoders

### v3

@@ 3.2 Blind Source Separation with Multi-Encoder Autoencoders
In this paper, we train a multi-encoder autoencoder for reconstructing underdetermined single-channel input mixtures where the encoders specialize in the feature subspaces of different sources as an emergent property of learning from the feature distribution of the data. Specifically, the encoders unmix the sources within the encoding space of the network, and the decoder remixes the sources such that each source component removed in the encoding space results in the source being removed from the output reconstruction. In 0(a), the training phase is shown during which all encoders $\bm{E}$ are active and the expected output of the network is a reconstruction $\hat{x}$ of the input signal $x$. However during inference, shown in 0(b), only one encoder is active and the expected output is the source signal $s^{n}$ that corresponds to the source feature subspace that the active encoder $E^{n}$ specializes in. In place of the encodings from non-active encoders, zero matrices of the same size are concatenated to the active encoder’s output encoding, keeping the position of the active encoding in place.
@@
where $\|B_{i,j}\|_{1}$ denotes the L1 norm of some block $B_{i,j}$. The sparse mixing loss is not applied to the final output layer of the network. In our standard formulation, the scalar $\alpha$ is an additional term equal to the reciprocal of the number of elements in each block $B_{i,j}$. However, for our experiments with PPG and ECG signals, another scaling scheme that assigns different penalties depending on each block’s position is devised in consideration of the linear and non-linear relationships between different sources in PPG or ECG signals, as is seen in Equation 21. The role of the proposed sparse mixing loss is illustrated in Figure 2. For the weight matrix $W$, the smallest squares have a value of one if they are white and a value of zero if they are black. In the input and output vectors, $v_{\text{in}}$ and $v_{\text{out}}$, the red, green, and blue colors represent three separate source encodings or mappings. In the case that all off-diagonal weights are exactly zero as illustrated by 1(a), no mixing occurs between the source encodings until the output layer (a convolutional layer with a kernel size of 1) of the network meaning that the signals are linearly mixed with the possibility of one non-linearity at the output [^1] For our experiments a sigmoid function is applied to the output for training with the binary cross entropy loss. If the triangles & circles mixtures did not include a distortion kernel after the sigmoid, the described scenario would be a valid solution.. However, for non-linearly mixed signals, off-diagonal blocks in each layer are expected to have at least some non-zero values as illustrated in 1(b).
@@
During inference, source separation is performed by leaving the $n^{\text{th}}$ encoder $E^{n}_{\theta_{n}}$ active and masking out all other encodings with zero vectors $\bm{0}$. The concatenation of the active encoding with the masked encodings $Z^{n}$ are passed into the decoder to give the source estimation $\hat{s}^{n}$ as illustrated in 0(b) and described in Algorithm 2 located in the appendices.
@@ 4.1 Triangles & Circles Toy Dataset
For this experiment, a fully convolutional multi-encoder autoencoder, detailed in 5(b) (located in Appendix B), is trained on the introduced toy dataset using the proposed method. Though there are only two sources for this dataset, three encoders are implemented to show the emergence of a dead encoder. Hyperparameter settings for the optimizer and all loss terms can be found in subsection C.1.
@@ 4.2 ECG and PPG Respiratory Source Separation
For both the ECG and PPG experiments, the same fully convolutional multi-encoder autoencoder, detailed in 5(a) (located in the appendices), is trained using the proposed method. Additionally, all hyperparameter settings are located in subsection D.1.
@@
In Figure 5, a simultaneously measured reference respiration signal, and the extracted respiration source signal using our method are visualized for both ECG and PPG experiments (more samples are in the supplementary material). In 4(a) the reference respiration signal shown is thoracic excursion, and for 4(b) the reference respiratory signal is nasal pressure. While the predicted respiratory sources are the result of the respiratory system’s functions they may also represent subtly different physiological processes than those measured by nasal flow or thoracic excursion (it may be more appropriate to describe these two signals as pseudo-references in this scenario so as not to confuse the common meaning of "reference" in source separation literature). Thus the predicted respiratory source and the reference respiration signal may have subtle or obvious differences in feature expression. For example, the predicted respiration source may have unique amplitude modulation patterns or appear flipped relative to the respiration reference as seen in 4(a). For these reasons, we cannot effectively evaluate the proposed method on this challenging problem with more standard metrics used in the source separation literature such as SAR, SIR, or SDR [94] (which are commonly used metrics for audio source separation tasks) because the ground truth respiratory source within a real-world ECG or PPG signal is not knowable and therefore what constitutes interference or artifacts is also unknowable. Even the reference signals have considerably different feature expressions from one another (as seen in the supplementary material). Instead, the proposed method is evaluated by extracting respiratory rate from the predicted source signals and for the two reference respiratory signals, nasal flow and thoracic excursion, which is achieved by measuring the periods between inhalation onsets or troughs for one-minute segments. Then the mean absolute error (MAE) between the respiratory rates the source estimation and each of the reference signals is calculated to determine the degree to which the predicted respiration sources preserve respiratory information relative to these references.
@@ B.3 Indentifiablility and Separability with Non-Linear Mixtures
7(a) shows a graph with three well-separated feature subspaces, $s^{a}$, $s^{b}$, and $s^{c}$, that correspond to the sources in a mixture. 7(b) shows a graph where the mixing system $\mathcal{F}$ maps the source feature subspaces into a new feature space where there may be increased ambiguity between the source feature subspaces or in the worst case complete ambiguity. The degree of ambiguity between sources within a mixture also determines how well sources can be recovered, if at all, which is referred to in the literature as separability or identifiability. The issue of separability in non-linear BSS has proven to be a challenging problem due to the additional ambiguities and indeterminacies often introduced by non-linear systems [23, 18]. The issue of separability in non-linear mixtures is unavoidable, and the method proposed in this paper aims to estimate sources even when indeterminacies are present as opposed to finding exact source reconstructions. Source estimations in this scenario may still provide useful representations as shown by our experimental results in subsection 4.2. In subsection 4.1, we further showed that the issues of identifiability and separability may be overcome through learning over a distribution of mixtures.

### v4

@@ 3.2 Blind Source Separation with Multi-Encoder Autoencoders
In this paper, we train a multi-encoder autoencoder for reconstructing underdetermined single-channel input mixtures where the encoders specialize in the feature subspaces of different sources as an emergent property of learning from the feature distribution of the data. Specifically, the encoders unmix the sources within the encoding space of the network, and the decoder remixes the sources such that each source component removed in the encoding space results in the source being removed from the output reconstruction. In 1(a), the training phase is shown during which all encoders $\bm{E}$ are active and the expected output of the network is a reconstruction $\hat{x}$ of the input signal $x$. However during inference, shown in 1(b), only one encoder is active and the expected output is the source signal $s^{n}$ that corresponds to the source feature subspace that the active encoder $E^{n}$ specializes in. In place of the encodings from non-active encoders, zero matrices of the same size are concatenated to the active encoder’s output encoding, keeping the position of the active encoding in place.
@@
where $\|B_{i,j}\|_{1}$ denotes the L1 norm of some block $B_{i,j}$. The sparse mixing loss is not applied to the final output layer of the network. In our standard formulation, the scalar $\alpha$ is an additional term equal to the reciprocal of the number of elements in each block $B_{i,j}$. However, for our experiments with PPG and ECG signals, another scaling scheme that assigns different penalties depending on each block’s position is devised in consideration of the linear and non-linear relationships between different sources in PPG or ECG signals, as is seen in Equation 14. The role of the proposed sparse mixing loss is illustrated in Figure 2. For the weight matrix $W$, the smallest squares have a value of one if they are white and a value of zero if they are black. In the input and output vectors, $v_{\text{in}}$ and $v_{\text{out}}$, the red, green, and blue colors represent three separate source encodings or mappings. In the case that all off-diagonal weights are exactly zero as illustrated by 2(a), no mixing occurs between the source encodings until the output layer (a convolutional layer with a kernel size of 1) of the network meaning that the signals are linearly mixed with the possibility of one non-linearity at the output [^1] For our experiments a sigmoid function is applied to the output for training with the binary cross entropy loss. If the triangles & circles mixtures did not include a distortion kernel after the sigmoid, the described scenario would be a valid solution.. However, for non-linearly mixed signals, off-diagonal blocks in each layer are expected to have at least some non-zero values as illustrated in 2(b).
@@
During inference, source separation is performed by leaving the $n^{\text{th}}$ encoder $E^{n}_{\theta_{n}}$ active and masking out all other encodings with zero vectors $\bm{0}$. The concatenation of the active encoding with the masked encodings $Z^{n}$ are passed into the decoder to give the source estimation $\hat{s}^{n}$ as illustrated in 1(b) and described in Algorithm 2 located in the appendices.
@@ 4.1 Triangles & Circles Toy Dataset
For this experiment, a fully convolutional multi-encoder autoencoder, detailed in 6(b) (located in Appendix B), is trained on the introduced toy dataset using the proposed method. Though there are only two sources for this dataset, three encoders are implemented to show the emergence of a dead encoder. Hyperparameter settings for the optimizer and all loss terms can be found in subsection C.1.
@@ 4.2 ECG and PPG Respiratory Source Separation
For both the ECG and PPG experiments, the same fully convolutional multi-encoder autoencoder, detailed in 6(a) (located in the appendices), is trained using the proposed method. Additionally, all hyperparameter settings are located in subsection D.1.
@@
In Figure 5, a simultaneously measured reference respiration signal, and the extracted respiration source signal using our method are visualized for both ECG and PPG experiments (more samples are in the supplementary material). In 5(a) the reference respiration signal shown is thoracic excursion, and for 5(b) the reference respiratory signal is nasal pressure. While the predicted respiratory sources are the result of the respiratory system’s functions they may also represent subtly different physiological processes than those measured by nasal flow or thoracic excursion (it may be more appropriate to describe these two signals as pseudo-references in this scenario so as not to confuse the common meaning of "reference" in source separation literature). Thus the predicted respiratory source and the reference respiration signal may have subtle or obvious differences in feature expression. For example, the predicted respiration source may have unique amplitude modulation patterns or appear flipped relative to the respiration reference as seen in 5(a). For these reasons, we cannot effectively evaluate the proposed method on this challenging problem with more standard metrics used in the source separation literature such as SAR, SIR, or SDR [94] (which are commonly used metrics for audio source separation tasks) because the ground truth respiratory source within a real-world ECG or PPG signal is not knowable and therefore what constitutes interference or artifacts is also unknowable. Even the reference signals have considerably different feature expressions from one another (as seen in the supplementary material). Instead, the proposed method is evaluated by extracting respiratory rate from the predicted source signals and for the two reference respiratory signals, nasal flow and thoracic excursion, which is achieved by measuring the periods between inhalation onsets or troughs for one-minute segments. Then the mean absolute error (MAE) between the respiratory rates the source estimation and each of the reference signals is calculated to determine the degree to which the predicted respiration sources preserve respiratory information relative to these references.
@@ B.3 Indentifiablility and Separability with Non-Linear Mixtures
8(a) shows a graph with three well-separated feature subspaces, $s^{a}$, $s^{b}$, and $s^{c}$, that correspond to the sources in a mixture. 8(b) shows a graph where the mixing system $\mathcal{F}$ maps the source feature subspaces into a new feature space where there may be increased ambiguity between the source feature subspaces or in the worst case complete ambiguity. The degree of ambiguity between sources within a mixture also determines how well sources can be recovered, if at all, which is referred to in the literature as separability or identifiability. The issue of separability in non-linear BSS has proven to be a challenging problem due to the additional ambiguities and indeterminacies often introduced by non-linear systems [23, 18]. The issue of separability in non-linear mixtures is unavoidable, and the method proposed in this paper aims to estimate sources even when indeterminacies are present as opposed to finding exact source reconstructions. Source estimations in this scenario may still provide useful representations as shown by our experimental results in subsection 4.2. In subsection 4.1, we further showed that the issues of identifiability and separability may be overcome through learning over a distribution of mixtures.